OpenStudy (anonymous):

Use the quadratic formula to solve the equation. If necessary, round to the nearest hundredth. x² + 3 = –4x

OpenStudy (unklerhaukus):

get your equation into this form \[ax^2+bx+c=0\]then find x with the quadratic formula\[x=\frac{-b± \sqrt{b^2-4ac}}{2a} \]

OpenStudy (anonymous):

remind me what the = symbol with the line is

OpenStudy (anonymous):

I am really bad at math :'(

OpenStudy (unklerhaukus):

± is + and - so there is two solution one when we use a + and the second when we use a -

OpenStudy (earthcitizen):

Firstly, re-arrange and\[x ^{2}+4x+3 =(x+1) (x+3) = 0\] factorise the equation and solve x= -1 0r -3. You don't need the quadratic formula. However, simply use \[a x^{2}+bx+c =0\] equate it a=1, b=4 and c=3 to the quadratic formula\[(x = (-b±\sqrt{(b^2-4 a c)})/(2 a))\] and solve